gRAPH THE SOLUTION of 5x+3y<15
5x-4y<20

Answer by mananth(16949)

5x+3y<15
5x-4y<20
First we find the boundary of each line
5x+3y=15............1
5x-4y=20............2
equation1
when x= 0 y= 5 ( 0 , 5 )
when y = 0 x= 3 ( 3 , 0 )
Equation 2
when x= 0 y= -5 ( 0 , -5 )
when y = 0 x= 4 ( 4 , 0 )
The Boundary

The Inequality


m.ananth@hotmail.ca